Find the Laplace transform, if it exists, of each of the following functions

(a) f(t) = eat

(b) f(t) = 1

(c) f(t) = t 


1
Expert's answer
2021-09-23T16:50:16-0400

(a)


"F(s)=L(f(t))=L(e^{at})=\displaystyle\int_{0}^{\infin}e^{at}e^{-st}dt"

"=-\dfrac{1}{s-a}\lim\limits_{A\to\infin}[e^{(a-s)t}]\begin{matrix}
 A \\
 0
\end{matrix}=-\dfrac{1}{s-a}(0-1)"




"=\dfrac{1}{s-a}"

(b)


"F(s)=L(f(t))=L(1)=\displaystyle\int_{0}^{\infin}(1)e^{-st}dt"




"=-\dfrac{1}{s}\lim\limits_{A\to\infin}[e^{-st}]\begin{matrix}
 A \\
 0
\end{matrix}=-\dfrac{1}{s}(0-1)"

"=\dfrac{1}{s}"

(c)


"F(s)=L(f(t))=L(t)=\displaystyle\int_{0}^{\infin}(t)e^{-st}dt"

"=-\dfrac{1}{s}\lim\limits_{A\to\infin}[te^{-st}+\dfrac{1}{s}e^{-st}]\begin{matrix}
 A \\
 0
\end{matrix}=-\dfrac{1}{s}(0-0+0-\dfrac{1}{s})"

"=\dfrac{1}{s^2}"